OpenStudy (anonymous):

Prove or disprove: There exist rational numbers a and b such that a^b is irrational. What kind of proof did you give?

OpenStudy (freckles):

For there exist proof you just come up with in a and b such that we you are trying to show is true! for example, Prove there exist integers m and n such that m+n=5 Proof: Set m=1 and n=4. Then m+n=1+4=5.\\ Proof done. Do you know what rationals are?

OpenStudy (freckles):

oh this was 2 days ago probably not around anymore :p

OpenStudy (loser66):

just "there exists", not all, hence arbitrary pick an irrational number, let say \(\sqrt 5\) then set a = 5, b = 1/2, hence \(a^b = \sqrt 5\) irrational!!